OpenStudy (anonymous):

Marley is a sales associate at a shoe store. Each week she earns $300 plus a 2% commission of all her sales. This week she earned $318. How much, in dollars, were Marley's shoe sales this week? (Do not put a $ in your answer.)

OpenStudy (jackellyn):

318-300=18 =commission this week x*0.02=18 x=18/0.02 x= what?

OpenStudy (anonymous):

x = 900 ?

OpenStudy (jackellyn):

exactly, good job @AshhSmith